Use Substitution ,ethod to find the simultaneous solution.
-3x+y=17
2x-3y= -23

Answer by robertb(5830)

y = 3x+17, the first equation. Put into the bottom equation:
2x - 3(3x + 17) = -23,
2x - 9x - 51 = -23,
-7x = 28,
x = -4. Then y = 3*-4 + 17 = -12 + 17 = 5. So y = 5.